OpenStudy (anonymous):

Solve this system of equations using the inverse of a matrix. -x + y = 8 4x + 3y = 12 x − 7y + z = 15 x = 4, y = 68, and z = 5.39 x = 68, y = 5.39, and z = 4 x = -1.76, y = 6.24, and z = 60.76 x = 6.24, y = -1.76, and z = 60.76 x = 72, y = -4, and z = 12.78 [Hint: AX = B and X = BA-1]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i surely would not use that method!

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

:)

OpenStudy (misty1212):

\[-x + y = 8\\ 4x + 3y = 12\] can be solved very easily

OpenStudy (anonymous):

how do you solve?

OpenStudy (misty1212):

multiply the first equation by \(-3\) and add

OpenStudy (misty1212):

or multiply the first equation by \(4\) and add

OpenStudy (misty1212):

either way

OpenStudy (anonymous):

-3x+-3y=-24?

OpenStudy (misty1212):

by \(-3\) you get \[3x-3y=-24\\ 4x+3y=12\] add and the y's go bye bye

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

once you have \(x\) and \(y\) plug them in to the third equation to find \(z\)

OpenStudy (anonymous):

so 3x=-24

OpenStudy (anonymous):

and 4x=12?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I am just going to take it as C